SYMMETRY-BASED TUNNELINGS IN HIGH-RESOLUTION ROVIBRATIONAL SPECTRA OF OCTAHE-DRAL MOLECULES

\underline{JUSTIN MITCHELL}, WILLIAM HARTER, Department of Physics, University of Arkansas, Fayetteville, AR 72701.

High-resolution spectra of spherical-top molecules are known to demonstrate rotational level clustering. This clustering is well described as a rotational phase-space effect^{a}. Multiple equivalent phase-space regions allow tunneling and thus splitting of the rotational clusters^{b}. So far this has been done with an ad hoc tunneling Hamiltonian. Similar splittings have been shown for low dimensional systems^{c}, also with an ad hoc parameterization. While ad hoc tunneling parameterization is simple to understand, it becomes extremely difficult to apply for higher symmetries and for locally low-symmetry clustering when many tunneling paths are possible. Symmetry-based parameterization mitigates this complication. This presentation will discuss how symmetry-based tunneling is applied for octahedral molecules and demonstrate how local-C_{1}, C_{2}, C_{3} or C_{4} clusters may be evaluated perturbatively. Connections to non-rotational systems, such as large amplitude motion, will be discussed as well.

aW.G. Harter and C.W. Patterson, Phys Rev Lett 38, 224 (1977)

bW.G. Harter and C.W. Patterson, J Chem Phys 66, 4872 (1977)

cJ.T. Hougen J Mol Spect 123, 197 (1987)